An instant oatmeal mix is considering adding flavors to its mix. 200 people tested the flavors and gave their preferences. Is there a preference for the flavor at the 0.05 level? State the hypotheses and identify the claim, find the critical value(s), compute the test value, make the decision, summarize the results.

Plain 20

Cinnamon 58

Apple 48

Maple 22

Peach 52
Answer #1

H0: Null Hypothesis: There is no preference for the flavor.

HA: Alternative Hypothesis: There is a preference for the flavor.

ndf = n - 1 = 5 - 1 = 4

= 0.05

From Table, critical value of = 9.4877

Expected Frequency for each flavor = 200/5 = 40

Test Statistic () is got as follows:

Observed (O) Expected (E) (O - E)2/E
20 40 10
58 40 8.1
48 40 1.6
22 40 8.1
52 40 3.6
Total = = 31.4

Since calculated value of = 31.4 is greater than critical value of = 9.4877, the difference is significant. Reject null hypothesis.

Conclusion:

The data support the claim that there is a preference for the flavor.